OpenStudy (anonymous):

Which value of x is the solution of 2x/5+1/3=7x-2/15? a) 3/5 b) 31/26 c) 3 d) 7

myininaya (myininaya):

multiply both sides by 15 5x+5=105x-2 -5x -5x 5=100x-2 +2 +2 7=100x - --- 100 100 x=7/100

myininaya (myininaya):

oops 6x+5=105x-2

myininaya (myininaya):

5=99x-2 7=99x x=7/99

OpenStudy (anonymous):

thanks hun:)